write a quaratic quation in standard form with -1/3 and 6 as its roots
Answer by nerdybill(7384)

Standard form would be aX^2 + bX + C = 0
if roots are -1/3 and 6
then factors are:
(x + 1/3) and (x - 6)
.
multiply the two to get the quadratic:
(x + 1/3)(x - 6) = 0
x^2 - 6x + x/3 - 2 = 0
x^2 - (18x)/3 + x/3 - 2 = 0
x^2 - (17/3)x - 2 = 0